OpenStudy (anonymous):

What is the x-intercept of the line with the equation ? -2x+1/2y=18 I think -9

OpenStudy (anonymous):

Medal & Fan Please Help

OpenStudy (haleyelizabeth2017):

Plug in 0 for y and solve for x :)

OpenStudy (haleyelizabeth2017):

Are there options?

OpenStudy (anonymous):

no & ok

OpenStudy (anonymous):

is -9 correct? @haleyelizabeth2017

OpenStudy (haleyelizabeth2017):

Give me a moment, sorry. I was helping another person lol

OpenStudy (anonymous):

your fine lol and ok

OpenStudy (haleyelizabeth2017):

\[-2x+\frac{1}{2}y=18\] Is this what it looks like?

OpenStudy (anonymous):

yes

OpenStudy (haleyelizabeth2017):

Okay, so: \[-2x+\frac{1}{2}(0)=18\]\[-2x=18\]\[x=-9\] You are correct!

OpenStudy (anonymous):

thank you can you help me with one more ?

OpenStudy (haleyelizabeth2017):

Sure. Tag me :)

OpenStudy (kobi):

−2x+12(0)=18 −2x=18 x=−9